OpenStudy (anonymous):

solve the equation... x^2 64 ---- = ---- x-8 x-8

OpenStudy (anonymous):

this is pretty simple because they have the same denominator, you know that the numerators must be equal. given that, x^2=64, and the square root of 64 is 8 OR -8. Now plug each of these into the equation to see if it works. for 8, you plug it in and have 8^2/(8-8)=64/(8-8) BUT 8-8 =0 and you cannot divide by 0. so you know 8 does not work. now testing with -8 (-8)^2/(-8-8)=64/(-8-8) 64/(-16)=64/(-16) which is TRUE. So your answer is -8.

OpenStudy (anonymous):

\[x= \pm8\]

jhonyy9 (jhonyy9):

x*2 /(x-8) =64/(x-8) --- x*2 =64 --- x*2 =8*2 --- x_1=64*(1/2) =8 and x_2 =64*(1/2) =-8